OpenStudy (anonymous):

What mass of water at 20.0°C must be allowed to come to thermal equilibrium with a 1.5-kg cube of aluminum initially at 163.0°C to lower the temperature of the aluminum to 57.0°C? Assume any water turned to steam subsequently recondenses.

OpenStudy (anonymous):

Still need help @biogirl?

OpenStudy (anonymous):

For thermal equilibrium: Qnet = 0 Qwater + QAl = 0

OpenStudy (anonymous):

Qwater = -QAl

OpenStudy (anonymous):

General equation: Q = m*C*\[DeltaT\]

OpenStudy (anonymous):

So... \[m _{w}*C _{w}*DeltaT _{w} = - m _{Al}*C _{Al}*DeltaT _{Al}\]

OpenStudy (anonymous):

Cw=4180 J/kg*K Tw=(57-20) mAl=1.5 kg Cal=904 J/kg*K TAl=(57-163) mw = ?

OpenStudy (anonymous):

One equation, one unknown, solve for the mass of the water.